Can you show the work on the following equation so I can "see how" it is done?
x^2 -6x -16 = 0
x^2 means x squared.
Thank you!!

Found 2 solutions by AKT, richard1234:
Answer by AKT(2)

x^2-6x-16=0
then
x^2-8x+2x-16=0
then
x(x-8)+2(x-8)=0,
(x-8) (x+2) =0,
therefor
x=8 or x=-2.
-2 is not a possible answer therefor x=8 is the final answer.

Answer by richard1234(7193)

This expression can be factored, because factors of -16 which add up to -6 are -8 and 2. Hence, it can be factored as

(x-8)(x+2) = 0. By the zero-property rule, x = 8 or x = -2.